MILLIMETER-WAVE SPECTROSCOPY OF THE VINYL RADICAL GENERATED BY UV LASER PHO-TOLYSIS IN A PULSED SUPERSONIC JET EXPANSION: DETERMINATION OF THE PROTON TUN-NELING SPLITTING

MASAAKI TOSHIMITSU, KENSUKE HARADA, \underline{KEIICHI TANAKA}, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

The rotational and proton tunneling spectra of the vinyl radical have been observed by millimeter-wave spectroscopy combined with a pulsed supersonic jet technique. The vinyl radical was generated by the 193 nm excimer laser photolysis of vinyl bromide. The pure rotational transitions, N_{K_{a}K_{c}} = 1_{01} − 0_{00}, 2_{02} − 1_{01}, 3_{03} − 2_{02}, and 4_{04} − 3_{03}, observed in the frequency region of 60 − 250 GHz, consisted of two components, 0^{+} and 0^{−}, confirming the proton tunneling motion of the vinyl radical. The observed rotational lines were split into fine and hyperfine components due to the spin-rotation interaction and the spin-nuclear spin interaction of the acetylenic (CH) as well as methylenic (CH_{2}) protons. The pure rotational spectra for K_{a} = 1 were also observed both for the 0^{+} and 0^{−} components. The rotatinal constants, spinrotation coupling constants, and hyperfine coupling constants for each tunneling components obtained were consistent with the results of the ESR spectroscopy in the Ar matrix^{a} and the infrared diode laser spectroscopy^{b}. The b-type Qbranch lines for the 0^{+} ← 0^{−} proton tunneling transition were observed around 200 GHz. The proton tunneling splitting was determined to be about 0.54 cm^{−1}, and the barrier height of the double minimum potential to be about 1250 cm^{−1}.

aP. H. Kasai, J. Am. Chem. Soc. 94, 5950 (1972)

bH. Kanamori, Y. Endo, and E. Hirota, J. Chem. Phys. 92, 197 (1990)